OpenStudy (anonymous):

Find the slope of the line 2x + 8y = 6 a) -1/4 b) 1/4 c) 4 d)-4

OpenStudy (anonymous):

can you put it in the Y = mX + b form?

OpenStudy (anonymous):

yes how???

OpenStudy (anonymous):

first, subtract 2x from both sides

OpenStudy (kira_yamato):

8y = 6 -2x y = 3/4 - x/4 (b)

OpenStudy (anonymous):

almost... but check your signs

OpenStudy (anonymous):

Y = -X/4 + 3/4

OpenStudy (kira_yamato):

Oh... Arigatou...

OpenStudy (kira_yamato):

Gomen ne, it should be (a)

OpenStudy (anonymous):

looks good to me... good job